OpenStudy (anonymous):

multiply and simplify by factoring. Assume that all expressions under radicals represent non-negative numbers 3sqrty^4 ^3sqrt81y^5

jimthompson5910 (jim_thompson5910):

\[\LARGE \sqrt[3]{y^4}\sqrt[3]{81y^5}\] \[\LARGE \sqrt[3]{y^4*81y^5}\] \[\LARGE \sqrt[3]{81y^9}\] \[\LARGE \sqrt[3]{27*3(y^3)^3}\] \[\LARGE \sqrt[3]{27}*\sqrt[3]{3}*\sqrt[3]{(y^3)^3}\] \[\LARGE 3*\sqrt[3]{3}*y^3\] \[\LARGE 3y^3\sqrt[3]{3}\] So \[\LARGE \sqrt[3]{y^4}\sqrt[3]{81y^5}=3y^3\sqrt[3]{3}\]